OpenStudy (anonymous):

What is the approximate value of cot pi/8?

OpenStudy (jkristia):

my approach would probably be to use the half angle formula for sin (x) and cos (x), since sin(pi/4) and cos(pi/4) are will defined.

OpenStudy (anonymous):

\[\cot (x)=\frac{ 1 }{ \tan(x) } \implies \cot \left( \frac{ \pi }{ 8 } \right)=\frac{ 1 }{ \tan\left(\frac{ \pi }{ 8 } \right) }\]

OpenStudy (anonymous):

@fprt Do you understand what I just did there? Can you compute the value now?

OpenStudy (anonymous):

Oh yeah that is the may I did it and got 145.9, but i wasn't sure if I did it incorrectly

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

okay so if it were in radians; would it be 145.9 pi rad/ 180 ??

OpenStudy (mertsj):

pi/4 is 45 degrees. pi/8 is 22.5 degrees. set your calculator in degrees. Type 1 divided by tan 22.5 equals. You should get 2.4

OpenStudy (anonymous):

This is funny. 1/(tan(pi/8)) shows up as 145.9 but 1/tan(22.5) shows up as 2.4? 0,o

OpenStudy (anonymous):

Oh because I'm not in degrees lol.

OpenStudy (anonymous):

Yes so, if you convert pi/8 radians to degrees, which is by multiplying it by 180/pi, you get 22.5 degrees. Then just do 1/tan(22.5) or you can do it directly as 1/tan(pi/8) if you're calculator is in radians mode. And the answer should be 2.4 like @Mertsj said. @fprt

OpenStudy (anonymous):

okay thanks y'all!